Grade 12th passMechanics

A grinding wheel is rotating at an angular speed of 8.6rad/s, when it is switched off.It comes to rest in 192s.Find
(a) Its angular acceleration
(b) The total angle turned through during the slow down.

Maha Ch
8 Years agoGrade 12th pass

1 Answer

vishal
ApprovedApproved Tutor Answer8 Years ago
angular acceleration=0.045 rad/s2(retardation)
total angle turned=821.76 degree
ang acc,a=w2-w1/t
w2=0 as after 192 s it comes to rest
t=192 s
so a=0-8.6/192=0.045 rad/s2
 
angle turned=w1t-1/2(a)t2=(8.6)192-1/2(0.045)(192)2=821.76 degree